zarkam21:

The conversion factor relating miles to meters is 1mile=1610m . The speed of light is 3.00× 10 8  m/s . How fast is this in miles per hour ( miles/h )?

zarkam21:

I got 10800000=10.8 x 10^5 miles/h

zarkam21:

@Vocaloid

Vocaloid:

hm not quite, a quick search tells me it should be in the 600,000,000 range so the first step would be to convert meters to miles by dividing by the conversion factor 1610m to miles (3e8)/1610 = 186335.403726 ish now that we have miles/second we just divide by 3600 since there are 3600 seconds in an hour

Vocaloid:

nonono, from (3e8)/1610 = 186335.403726 meters/second we ~divide~ by 3600 to get to seconds whoops

zarkam21:

51.75983437

Vocaloid:

** meant to say multiply

Vocaloid:

so 186335.403726 * 3600

zarkam21:

670807453.4136

zarkam21:

Wait :S

Vocaloid:

yeah that should be it w/ sig figs it ends up being like 6.71 * 10^8 mph

zarkam21:

The conversion factor relating feet to meters is 1ft=0.305m . Keep in mind that when using conversion factors, you want to make sure that like units cancel leaving you with the units you need. You have been told that a certain house is 164 m 2 in area. How much is this in square feet?

zarkam21:

164 m^2(1ft/0.305m)=164^2*0.305=8203.28

Vocaloid:

this is a good question because square units require squared conversion factors if 1ft=0.305m then 1ft^2 = 0.305m^2 so 164 m^2 * (1ft^2/ 0.305m^2) should do it

Vocaloid:

164 m^2 * (1ft^2/ 0.305^2 m^2) should do it

zarkam21:

1762.966944

zarkam21:

so 1763

zarkam21:

@Vocaloid

zarkam21:

@jhonyy9

jhonyy9:

what is your question ?

jhonyy9:

the speed of light is 3×10^8 m/s do you know this ?

zarkam21:

@Vocaloid

Vocaloid:

hm, close, since we are only given 3 sig figs then it's 1760

zarkam21:

COuld you explain what I did wrong?

Vocaloid:

did the program say it was wrong? that's weird

zarkam21:

No no lol its right . I just wante dto see where I made my mistake

Vocaloid:

oh this part 164 m^2 * (1ft^2/ 0.305^2 m^2) = 1762.966944 is good since the number given in the problem is 164, that only has 3 sig figs so we round 1762.9 whatever down to 1760

zarkam21:

oh okay so we would go down vs. go up to 1763?

Vocaloid:

1763 has 4 sig figs the rules of rounding say that the 6 has to be rounded down because 3 < 5 so we round down to 1760

zarkam21:

okay understood. thanks